1.078 g sample of a compound containing only carbon, hydrogen, and oxygen is burned in an excess of dioxygen, producing 2.451 g CO2 and 1.003 g H2O.

1
Expert's answer
2022-06-03T05:54:03-0400

First we calculate the masses of carbon and hydrogen from their oxide (CO2 and H2O).

Mass of C = 2.451 g CO2 × 12 g C/ 44 g CO2

                = 0.668 g

Mass of H = 1.003 g H2O × 2 g H/ 18 g H2O

                = 0.111 g

Mass of C + Mass of H = 0.779 g

This is less than the mass of sample. The missing mass must be caused by O.

Mass of O = (1.078 - 0.779)g = 0.299 g


mol of C = 0.668 g / 12 g/mol = 0.056

mol of H = 0.111 g / 1 g/mol = 0.111

mol of O = 0.299 g / 16 g/mol = 0.019


C : H : O = 0.056 : 0.111 : 0.019 = 3 : 6 : 1

The empirical formula of the compound is C3H6O